Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 433     Entry time: 03 Nov 2023, 02:08
Author: Johannes Moeller 
Group: HED 
Subject: Shift summary 
  • stable 18 keV SASE delivery
  • measurement of siemens star in the morning
  • alignment of phase plate for the nano focussing lenses
  • testing and debugging of the pulse on demand macro
  • holography test measurements and optimization of online analysis
  • alignment of user samples

Problems:

  • FSSS ScannerX had to be shutdown and initiated quite often
  • We lost the output of the pp laser at 22:30. Laser OCD could solve the problem and the laser was back at 23:45.
ELOG V3.1.4-7c3fd00